darkknight:

Calc help on Related Rates and Implicit Differentiation problems Question 1

darkknight:

darkknight:

So what is the formula for the volume of this? Is is \[v=1/3*r^2*h*\pi\] For part A i did \[1/(b-a) * \int\limits_{a}^{b}r *dr = \] \[1/(103/20-3/20)*\int\limits_{3/20}^{103/20}* 1/20(3+h^2)dh = .605\]

darkknight:

I got 103/20 from plugging in h =10 into the formula that it gives and got 3/20 from plugging in h=0

darkknight:

@dude @tranquility

Mathery:

Let's try to label that funnel first.

darkknight:

?

darkknight:

@florisalreadytaken

Florisalreadytaken:

you are looking for b right? what you did above, i dont get its logic -- youve overcomplicated it, and in a wrong way also

darkknight:

so \[r=1/20(3+h^2)\] so to get the average value for the radius, don't we plug in 0 and 10 in for h (because 0< h< 10 INClusive) into the formula and get "a" and "b" ?

Florisalreadytaken:

no b is 10 a is 0

darkknight:

Oh, i see now, let me redo part a

darkknight:

So what is the formula for the volume of this? Is is \[v=1/3*r^2*h*\pi\] For part A i did \[1/(b-a) * \int\limits_{a}^{b}r *dr = \] \[1/(10-0)*\int\limits_{0}^{10}* 1/20(3+h^2)dh = 109/60=1.816666\] *** changed it ****

Florisalreadytaken:

omg i was doing b and i was like what is this guy doing... and then i looked at the image again and i realised youre doing A -- its 2am so yeah do not expect much that looks all good

darkknight:

oh alr, so the first was correct or the second for a?

Florisalreadytaken:

avg value is ok have you done the rest?

Florisalreadytaken:

for b use \[ \pi \int\limits_{a}^{b}r^2 *dr \]

darkknight:

Not quite, so for b to find the volume of the funnel I have to write an eqn for the volume in terms of r and h. But would the eqn be v = 1/3*pi *r^2*h or what Not sure what the volume formula for a funnel is.......

darkknight:

florisalreadytaken wrote:
for b use \[ \pi \int\limits_{a}^{b}r^2 *dr \]
why do we use this, I don't really get it if you could explain

Florisalreadytaken:

well yeah you are right its quite a complicated figure you want to do the same you did with your previous post right? separate it into circles do you get he idea of it?

Florisalreadytaken:

treat it like a cylinder

Florisalreadytaken:

One last thing i want to point out be4 i go haha I meant dh not dr That would be the height I just copied your latex thats why its dr lol : )

darkknight:

How to do part c though? @az maybe if you have time

AZ:

@darkknight wrote:
So what is the formula for the volume of this? Is is \[v=1/3*r^2*h*\pi\] For part A i did \[1/(b-a) * \int\limits_{a}^{b}r *dr = \] \[1/(10-0)*\int\limits_{0}^{10}* 1/20(3+h^2)dh = 109/60=1.816666\] *** changed it ****
I got kind of confused reading all the replies so I just want to clarify that this is correct for part A and then Flor told you how to find the volume of the funnel (and it's dh not dr as he later corrected the typo)

darkknight:

florisalreadytaken wrote:
for b use \[ \pi \int\limits_{a}^{b}r^2 *dr \]
alr cool, I just don't get where ^^^ comes from like why is tat the eqn for the volume and also confused on part c

AZ:

The volume of the funnel is going to be the sum of the areas of each consecutive circle that's formed as a cross section at different heights so \[ \text{Volume} = \pi \int\limits_{a}^{b}r^2 ~dh \] Volume = pi * r^2 h \(\pi\) is a constant so it's taken out and then you're integrating it with respect to the height so dh

AZ:

For the related rates part c, we know r = 1/20 (3 + h^2) you can distribute the 1/20 inside and we get \( r = \dfrac{3}{20} + \dfrac{h^2}{20}\) now it's easier to differentiate and you won't make any silly mistakes We want to know the rate at which the height is decreasing with respect to time (that is dh/dt) we know r = whatever can you differentiate that with respect to time so what is dr/dt = ?? and then you have to use what the question states, they tell us that the radius of the liquid is decreasing at a rate of 1/5 inch per second so that means you substitute -1/5 in place of dr/dt and then all you have to do is solve for dh/dt and that's your final answer

darkknight:

az wrote:
The volume of the funnel is going to be the sum of the areas of each consecutive circle that's formed as a cross section at different heights so \[ \text{Volume} = \pi \int\limits_{a}^{b}r^2 ~dh \] Volume = pi * r^2 h \(\pi\) is a constant so it's taken out and then you're integrating it with respect to the height so dh
I was afk and thinking about the problem and just realized why this is, ty

AZ:

@az wrote:
For the related rates part c, we know r = 1/20 (3 + h^2) you can distribute the 1/20 inside and we get \( r = \dfrac{3}{20} + \dfrac{h^2}{20}\) now it's easier to differentiate and you won't make any silly mistakes We want to know the rate at which the height is decreasing with respect to time (that is dh/dt) we know r = whatever can you differentiate that with respect to time so what is dr/dt = ?? and then you have to use what the question states, they tell us that the radius of the liquid is decreasing at a rate of 1/5 inch per second so that means you substitute -1/5 in place of dr/dt and then all you have to do is solve for dh/dt and that's your final answer
forgot to mention, you would also plug in h = 3 along with dr/dt = -1/5 into what you get when you differentiate \( r = \dfrac{3}{20} + \dfrac{h^2}{20}\) with respect to 'time' or 't' because they're asking at h = 3, dr/dt is -1/5 so what is dh/dt equal to?

darkknight:

\[r=3/20+h^2/20\] \[dr/dt=0+h/10*dh/dt\] -1.5 = 0 + 3/10 * dh/dt -1.5*10/3 = dh/dt

AZ:

\(\dfrac{\text{dr}}{\text{dt}} = \dfrac{d}{\text{dt}} \left(\dfrac{3}{20} + \dfrac{h^2}{20}\right) = 0 + \left( \dfrac{1}{20} \cdot\dfrac{d}{\text{dt}} \left(h^2\right) \right) \) \( = \dfrac{1}{20} \cdot 2h\cdot \dfrac{\text{dh}}{\text{dt}} \) \(\dfrac{\text{dr}}{\text{dt}} = \dfrac{h}{10} \cdot \dfrac{\text{dh}}{\text{dt}}\) so, yup good job but be careful, it's -1/5 not -1.5 but other than that, you're on the right track now :)

darkknight:

oh alr, yeah My head, I "mis-saw" that lol. Thanks (:

AZ:

You're welcome :)